# Plunge spectra as discriminators of black hole mimickers

### v1

@@ Abstract
As the compactness of these black hole mimickers are similar to that of black holes, they will reflect GW perturbations on a surface very close to the location of the event horizon of a black hole of comparable mass and spin. In the case of a Schwarzschild-like black hole mimicker, the reflection of perturbations takes place on a surface located very close to $2M$, say, $r_{s}=2M(1+\epsilon)$ such that $\epsilon\ll 1$, even as small as $\mathcal{O}(l_{p}/M)$. The perturbation at each frequency will be reflected completely or partially, depending on the internal properties of the system. Modifying the boundary condition imposed on perturbations near $2M$ with a black hole exterior is expected to affect the system’s response and behaviour, which has pointed towards some key observational tools which could be used to identify such black hole mimickers, this includes their quasinormal mode (QNM) spectrum Maggio:2020jml ; Laghi:2020rgl ; Saketh:2024ojw , tidal deformability Cardoso:2017cfl ; Nair:2022xfm ; Chakraborty:2023zed ; Berti:2024moe ; Silvestrini:2025lbe and more Cardoso:2017cqb ; Mark:2017dnq ; LongoMicchi:2019wsh ; Datta:2019epe ; Datta:2020rvo ; Agullo:2020hxe ; Chakraborty:2022zlq ; Chakravarti:2023wlc ; Datta:2024vll ; Bambi:2025wjx .
@@
Such orders of magnitude deviation at higher frequencies, accompanied by the sheer volume of individual observations that could be made possible with next-generation detectors such as LISA Babak:2017tow , compounded by the recent developments in using stacking methods Yang:2017zxs to enhance the signal-to-noise ratio (SNR) of GW events, suggests that despite the current difficulties with the SNR in the plunge phase of extreme mass ratio (EMR) events, the GW radiation during plunge could potentially become a powerful tool to identify black hole mimickers.
@@ II.2.2 New avenues at higher frequencies
Let us begin by noticing that for the $(\ell,m)=(2,2)$ modes of the metric perturbations in Equation 1, something interesting happens when $M\omega\gtrsim 0.39$. Above this, we will have the effective potential for the Zerilli modes to be less than $\omega^{2}$, meaning that the perturbing frequencies can directly penetrate the near-horizon. This observation suggests that for larger values of $\omega$, $\omega\gtrsim\omega_{\mathrm{th}}=0.39/M$, we can expect a quantitative change in the energy spectrum for the perturbation. In what follows, we will describe what to expect from a black hole mimicker when $M\omega\gtrsim 0.39$.
@@
In the analysis thus far, we have effectively set the source term in the Fourier domain to be equal to zero when $r_{p}<r_{s}$, and when $r_{p}>r_{s}$, we are setting it to $S_{\ell m\omega}^{(\pm)}(r)$. This amounts to stopping the time domain analysis at $t_{s}=t_{p}(r_{s})$; it is known that this could result in corrections in the frequency domain McKechan:2010kp ; Chen:2024ieh as in real events the source term will not vanish post $t_{s}$. This is not a concern in the case of a plunge into a black hole, as here the point particle asymptotes to $2M$ in infinite observer time; while in the context of a black hole mimicker, the time $t_{s}$ is finite. We will elaborate on this in the upcoming sections.
@@
We observed that just like the $dE/d\omega$ associated with the radial infall as given in Sec. III and Sec. III; for the geodesic plunge with non-zero angular momentum, we regain the characteristic resonances at lower frequencies and orders of magnitude deviations above $\omega_{\mathrm{th}}$. We observe that for sufficiently large values of $\omega$, the spectrum displays a behaviour similar to Equation 18. However, from Sec. III we can note that the falloff for different plunge orbits are different as it changes with the specific angular momentum, $\mathcal{L}$ of the point particle, however we observe that the difference is greatly suppressed in relation to the black hole case and the tail is ‘almost universal’ [^2] We thank Sayak Datta for pointing out that the high-frequency behaviour appears to be less sensitive to $\mathcal{L}$..
@@
As alluded to in the earlier section, the analysis of the plunge, which we have carried out in the frequency domain, maps to stopping the time analysis at $t_{p}(r_{s})$. This could result in corrections in the frequency domain McKechan:2010kp ; Chen:2024ieh , in relation to real observations where we can expect extra energy at each frequency bin due to the non-vanishing source term post $t_{s}$. These correction in the $M\omega<1$ range will be of the form $\Sigma_{i}c_{i}\omega^{i}$, with $c_{i}$’s that are determined by the full source term of the specific black hole mimicker (see Appendix B). However, by comparing the results of our analysis with those of greybody factors presented in Rosato:2025byu , we conclude that our work captures the correct features of the energy spectrum for the plunge into a black hole mimicker, in the presented frequency range.
We also note that the energy spectrum of the GW radiation emitted during a plunge can be expected to be relatively stable against environmental perturbations, as the grey body factor itself has been shown to be stable against environmental effects Oshita:2024fzf ; Rosato:2024arw . Thus, the energy spectrum of the GW radiation during plunge can be expected to be a smoking gun test for potential black hole mimickers.
@@ V Conclusion
We noted that the plunge could serve as a direct probe for near-horizon features as it excites GW perturbations at frequencies where the effective potential is less than the square of the perturbing frequency. By treating the direct plunge (both radial and non-radial) as a surrogate for the GUI plunge from the ISCO, we studied the GW spectrum at low as well as high frequencies.
@@ Acknowledgements
Let us assume that the observed black hole mimicker has a source term that includes information about what happens after the secondary reaches $r_{s}$, we can express it in the time domain as $\mathcal{S}_{\ell m}(t)$. In our model, we restricted ourselves to the plunge, and thus we have set the source term post $t_{s}=t_{p}(r_{s})$ to zero. So what we have done is our analysis of the plunge is set $S_{\ell m}^{(\pm)}(t)\approx\Theta(t_{c}-t)\mathcal{S}_{\ell m}(t)$, with $S^{(\pm)}_{\ell m}(t)$ being the plunge source term (as in Equation 4 and Equation 16) up to $t_{c}$ .
We wish to quantify how much the full source term for the black hole mimicker in the Fourier domain $\mathcal{S}_{\ell m\omega}=\int_{-\infty}^{\infty}\mathcal{S}_{\ell m}(t)e^{i\omega t}\,dt$ deviates from the model source therm, which we have used $S_{\ell m\omega}=\int_{-\infty}^{t_{c}}{S}_{\ell m}(t)e^{i\omega t}\,dt$. So we define the deviation $\Delta_{\ell m}(\omega)$ as
@@
$=-\!\int_{t_{c}}^{\infty}\mathcal{S}_{\ell m}(t)e^{i\omega t}\,dt~.$
@@
$\Delta(\omega)=-\sum_{n=0}^{\infty}\frac{(i\omega)^{n}}{n!}\,\mu_{n},~\mu_{n}=\int_{t_{c}}^{\infty}t^{n}\mathcal{S}_{\ell m}(t)\,dt.$ (26)
@@
One could also show that such a hard cut-off also yields a correction at $M\omega\gg 1$ of the form
$\Delta_{\ell m}(\omega)=e^{i\omega t_{c}}\!\left[\frac{\mathcal{S}_{\ell m}(t_{c})}{i\omega}-\frac{\mathcal{S}^{\prime}_{\ell m}(t_{c})}{(i\omega)^{2}}+\cdots\right].$ (27)

### v2

@@ Abstract
As the compactness of these black hole mimickers are similar to that of black holes, they will reflect GW perturbations on a surface very close to the location of the event horizon of a black hole of comparable mass and spin. In the case of a Schwarzschild-like black hole mimicker, the reflection of perturbations takes place on a surface located very close to $2M$, say, $r_{s}=2M(1+\epsilon)$ such that $\epsilon\ll 1$, even as small as $\mathcal{O}(l_{p}/M)$. The perturbation at each frequency will be reflected completely or partially, depending on the internal properties of the system. Modifying the boundary condition imposed on perturbations near $2M$ with a black hole exterior is expected to affect the system’s response and behaviour, which has pointed towards some key observational tools which could be used to identify such black hole mimickers, this includes their quasinormal mode (QNM) spectrum Maggio:2020jml ; Laghi:2020rgl ; Saketh:2024ojw , tidal deformability Cardoso:2017cfl ; Nair:2022xfm ; Chakraborty:2023zed ; Berti:2024moe ; Silvestrini:2025lbe and more Cardoso:2017cqb ; Mark:2017dnq ; LongoMicchi:2019wsh ; Datta:2019epe ; Datta:2020rvo ; Agullo:2020hxe ; Chakraborty:2022zlq ; Chakravarti:2023wlc ; Datta:2024vll ; Bhattacharya:2025xko ; Bambi:2025wjx .
@@
Such orders of magnitude deviation at higher frequencies, accompanied by the sheer volume of individual observations that could be made possible with next-generation detectors such as LISA Babak:2017tow , compounded by the recent developments in using stacking methods Yang:2017zxs to enhance the signal-to-noise ratio (SNR) of GW events, suggests that despite the current difficulties with the SNR in the plunge phase of extreme mass ratio (EMR) events, the GW radiation during plunge could plausibly become a powerful tool to identify black hole mimickers over a four years observation run of LISA.
@@ II.2.2 New avenues at higher frequencies
Let us begin by noticing that for the $(\ell,m)=(2,2)$ modes of the metric perturbations in Equation 1, something interesting happens when $M\omega\gtrsim 0.39$. Above this, we will have the effective potential for the Zerilli modes to be less than $\omega^{2}$, meaning the waves can transmit more effectively through the potential barrier and probe the near-horizon region. This observation suggests that for larger values of $\omega$, $\omega\gtrsim\omega_{\mathrm{th}}=0.39/M$, we can expect a quantitative change in the energy spectrum for the perturbation. In what follows, we will describe what to expect from a black hole mimicker when $M\omega\gtrsim 0.39$.
@@
In the analysis thus far, we have effectively set the source term in the Fourier domain to be equal to zero when $r_{p}<r_{s}$, and when $r_{p}>r_{s}$, we are setting it to $S_{\ell m\omega}^{(\pm)}(r)$. This amounts to stopping the time domain analysis at $t_{s}=t_{p}(r_{s})$; it is known that this could result in corrections in the frequency domain McKechan:2010kp ; Chen:2024ieh as in real events the source term will not vanish post-$t_{s}$. This is not a concern in the case of a plunge into a black hole, as here the point particle asymptotes to $2M$ in infinite observer time; while in the context of a black hole mimicker, the time $t_{s}$ is finite. We will elaborate on this in the upcoming sections.
@@
We observed that just like the $dE/d\omega$ associated with the radial infall as given in Sec. III and Sec. III; for the geodesic plunge with non-zero angular momentum, we regain the characteristic resonances at lower frequencies and orders of magnitude deviations above $\omega_{\mathrm{th}}$. We observe that for sufficiently large values of $\omega$, the spectrum displays a behaviour similar to Equation 18.
However, from Sec. III we can note that the falloff for different plunge orbits are different as it changes with the specific angular momentum, $\mathcal{L}$ of the point particle, however we observe that the difference is greatly suppressed in relation to the black hole case and the tail is ‘almost universal’ [^2] We thank Sayak Datta for pointing out that the high-frequency behaviour appears to be less sensitive to $\mathcal{L}$.. Similar universality has also been reported in Oshita:2023cjz ; Okabayashi:2024qbz , where the high-frequency GW spectrum becomes largely insensitive to the detailed properties of the source, even in the case of rotating black holes. This universality of the tail against the exact details of the source term further strengthens the case for using a direct plunge as a surrogate for the post-inspiral GUI plunge in studying high-frequency deviations. While the analytic arguments presented in Sec. II.2.1 justifies it at low frequencies.
@@
As alluded to in the earlier section, the analysis of the plunge, which we have carried out in the frequency domain, maps to stopping the time domain analysis at $t_{p}(r_{s})$. This could result in corrections in the frequency domain McKechan:2010kp ; Chen:2024ieh , in relation to real observations where we can expect the energy at each frequency bin to change due to the non-vanishing source term post-$t_{s}$. However, we note that for the range of frequencies explored in this paper, truncating the source at $t_{s}$ does not significantly alter the low frequency resonances as well as the spectral excess beyond $\omega_{th}$ (see Appendix B). Further, by comparing the results of our analysis with those of greybody factors presented in Rosato:2025byu , we conclude that our work captures the correct features of the energy spectrum for the plunge into a black hole mimicker, in the presented frequency range.
We note that while black hole spectroscopy based on discrete ringdown QNMs is a powerful probe of potential black hole mimickers, the plunge energy spectrum $dE/d\omega$ carries complementary physical information that may be lost in standard QNM-based ringdown fits. In practice, ringdown analyses model the late-time signal as a superposition of a few damped sinusoids and fit the data in a time window, compressing the response into a small set of complex frequencies and excitation amplitudes, where the fundamental mode dominates. By contrast, the energy spectrum explicitly resolves the full frequency-dependent response of the system to the plunge: it simultaneously captures the low-frequency resonance comb associated with the mimicker QNM spectrum and the qualitative break with an enhanced high-frequency tail for $\omega\gtrsim\omega_{th}$. As illustrated in Sec. II.2.2, this tail represents a continuous scattering process where waves transmit through the photon sphere barrier, reflect off the mimicker surface, and scatter back to infinity; this non-resonant mechanism cannot be efficiently parametrised by a sum of QNMs. This tail instead is related to the graybody factors (see Appendix C) of the black hole mimicker Oshita:2023cjz ; Okabayashi:2024qbz ; Rosato:2025byu , whose high-frequency behaviour has recently been shown to be stable under small environmental perturbations Oshita:2024fzf ; Rosato:2024arw . Thus, the energy spectrum of the GW radiation during plunge can be expected to be a smoking gun observable to identify black hole mimickers.
To assess the observational prospects of these plunge spectral features with LISA, let us note that individual EMRI plunges are expected to contribute only a small fraction of the total accumulated SNR, so resolving the resonance comb and the enhanced high-frequency tail in a single event may be challenging. As a rough scaling estimate, for a typical LISA EMRI with an SNR of $\rho_{\rm total}\sim 20$ accumulated over $\sim 10^{4}\text{--}10^{5}$ cycles Babak:2017tow . If only a few cycles contribute in the plunge/transition band, a naive duration-based estimate suggests an effective plunge SNR of at most $\rho_{\rm plunge}\sim\mathcal{O}(10^{-1})$}, with the precise value depending on the transition duration, the source parameters, and the LISA noise curves.
However, we note that the spectral features identified for these EMRIs are controlled entirely by the properties of the primary. This opens the possibility of combining information from multiple distinct EMRIs by employing the long inspiral preceding the plunge to determine the plunge time and the central-object parameters well enough to enable coherent alignment of the signals. With coherent alignment, the effective SNR can be expected to improve by $\propto\sqrt{N}$, analogous to Yang:2017zxs . If we aim to stack such signals to an effective plunge SNR of $\rho_{*}\sim 5$. It would require us to coherently stack $N\sim(\rho_{*}/\rho_{\rm plunge})^{2}\sim 10^{3}\text{--}10^{4}$ events. Population studies indicate that LISA may detect anywhere between a few and a few thousand EMRIs per year, subject to astrophysical conditions Babak:2017tow , so accumulating sufficient events over a four-year mission duration is plausible.
We further note that the high-frequency excess above $\omega_{th}$ is a generic feature of black hole mimickers with nonzero reflectivity $R$. For mimickers with detector frame masses in the LISA EMRI range ($M_{z}\sim 10^{5}\!-\!10^{7}M_{\odot}$), the threshold frequency $\omega_{th}$ lies within the LISA band, making this feature a particularly promising observational target with such stacking strategies. A quantitative assessment of waveform systematics and stacking strategies for LISA is left to future work.
@@ V Conclusion
We noted that the plunge could serve as a direct probe for near-horizon features as it excites GW perturbations at frequencies where the effective potential is less than the square of the perturbing frequency. Since the GW radiation during a direct plunge is not monochromatic Sago:2002fe ; Silva:2023cer , just like the post-ISCO GUI plunge in an EMRI Hadar:2009ip ; Rom:2022uvv . By treating the direct plunge (both radial and non-radial) as a surrogate for the GUI plunge from the ISCO, we studied the GW spectrum at low as well as high frequencies.
@@ Acknowledgements
Let us assume that the observed black hole mimicker has a source term that includes information about what happens after the secondary reaches $r_{s}$, we can express it in the time domain as $\mathcal{S}_{\ell m}(t)$. In our model, we restricted ourselves to the plunge, and thus we have set the source term post $t_{s}=t_{p}(r_{s})$ to zero. So what we have done is our analysis of the plunge is set $S_{\ell m}^{(\pm)}(t)\approx\Theta(t_{s}-t)\mathcal{S}_{\ell m}(t)$, with $S^{(\pm)}_{\ell m}(t)$ being the plunge source term (as in Equation 4 and Equation 16) up to $t_{s}$ .
We wish to quantify how much the full source term for the black hole mimicker in the Fourier domain $\mathcal{S}_{\ell m\omega}=\int_{-\infty}^{\infty}\mathcal{S}_{\ell m}(t)e^{i\omega t}\,dt$ deviates from the model source therm, which we have used $S_{\ell m\omega}=\int_{-\infty}^{t_{s}}{S}_{\ell m}(t)e^{i\omega t}\,dt$. So we define the deviation $\Delta_{\ell m}(\omega)$ as
@@
$=-\!\int_{t_{s}}^{\infty}\mathcal{S}_{\ell m}(t)e^{i\omega t}\,dt~.$
@@
$\Delta(\omega)=-\sum_{n=0}^{\infty}\frac{(i\omega)^{n}}{n!}\,\mu_{n},~\mu_{n}=\int_{t_{s}}^{\infty}t^{n}\mathcal{S}_{\ell m}(t)\,dt.$ (26)
@@
One could also show that such a hard cut-off also yields a correction at $M\omega\gg 1$ (Which is much larger than the largest value of $M\omega$ considered in this work.) of the form
$\Delta_{\ell m}(\omega)=e^{i\omega t_{s}}\!\left[\frac{\mathcal{S}_{\ell m}(t_{s})}{i\omega}-\frac{\mathcal{S}^{\prime}_{\ell m}(t_{s})}{(i\omega)^{2}}+\cdots\right].$ (27)
@@
To assess whether the truncation at $t_{s}$ could meaningfully affect the spectral excess beyond $\omega_{th}$ discussed in the main text, we next provide an order of magnitude estimate of the induced correction at a representative point, $M\omega\simeq 0.7>\omega_{th}$.
Order of magnitude estimate of truncation-induced corrections at a representative frequency
An estimate of the order of magnitude contribution to the observed GW energy spectrum from the omitted Fourier contribution $\Delta_{\ell m}(\omega)$ follows from the expectation that the post-$t_{s}$ source persists over a characteristic duration $T_{\rm post}$ of order the primary scale, i.e. $T_{\rm post}=\mathcal{O}(M)$. Further since the time domain source $\mathcal{S}_{\ell m}(t)$ is sourced by a body of mass $\mu$, we can expect $\mathcal{S}_{\ell m}(t)\sim\mu$, we can therefore parameterise
$|\Delta_{\ell m}(\omega)|\sim\eta\,\mu\,T_{\rm post},\qquad\eta=\mathcal{O}(1),$ (28)
where $\eta$ absorbs the dependence on the detailed post-$t_{s}$ source profile.
To estimate the change in the energy spectrum by incorporating $\Delta_{\ell m}(\omega)$, we can start by noticing that the RW/Zerilli equations are linear. As a result, the outgoing frequency-domain amplitudes are linear functionals of the source (cf. Equation 12). So the truncation induces a shift
$C^{(\pm){\rm,out}}_{\ell m\omega}\ \longrightarrow\ C^{(\pm){\rm,out}}_{\ell m\omega}+\delta C^{(\pm){\rm,out}}_{\ell m\omega},\,\delta C^{(\pm){\rm,out}}_{\ell m\omega}\sim\kappa\,\Delta_{\ell m}(\omega),$ (29)
with $\kappa=\mathcal{O}(1)$, encoding the effects from the integral over the homogeneous solution.
For a given mode, RW/Zerilli( or the corresponding effective amplitude combination entering Equation 11), the spectrum is quadratic in the complex amplitude,
$\frac{dE}{d\omega}=\mathcal{K}(\omega)\,|C^{(\pm){\rm out}}_{\ell m\omega}|^{2},$ (30)
where $\mathcal{K}(\omega)$ is the known real prefactor as infered from Equation 11. Hence the spectrum after incorporating $\Delta_{\ell m}(\omega)$ is
[…]
$=\frac{1}{\mu^{2}}\frac{dE}{d\omega}\left(1+2\alpha\cos\phi+\alpha^{2}\right),$
where we introduced
$\alpha\equiv\frac{|\delta C^{(\pm){\rm out}}_{\ell m\omega}|}{|C^{(\pm){\rm out}}_{\ell m\omega}|},\qquad\phi\equiv\arg(\delta C^{(\pm){\rm out}}_{\ell m\omega})-\arg(C^{(\pm){\rm out}}_{\ell m\omega}).$ (32)
Equation (31) makes the leading truncation effect explicit and shows it to be an interference correction ($\propto\alpha\cos\phi$), whose sign is controlled by the unknown relative phase $\phi$ of the omitted contribution.
Based on the above discussion, we can now estimate the corrected energy flux at a representative point, say $M\omega\simeq 0.7$, from Sec. III. For the dominant $(\ell,m)=(2,2)$ contribution, we define the effective amplitude $\mathcal{C}_{22\omega}$ through Equation 11 so that
$\frac{dE_{22}}{d\omega}=\frac{3\omega^{2}}{8\pi^{2}}\,\mathcal{C}_{22\omega}^{2}\quad\Rightarrow\quad\frac{\mathcal{C}_{22\omega}}{\mu}=\sqrt{\frac{8\pi^{2}}{3\omega^{2}}\left(\frac{1}{\mu^{2}}\frac{dE_{22}}{d\omega}\right)},$ (33)
Fig. 4: Plot of the graybody factor against $M\omega$ for the $(\ell,m)=(2,2)$ modes of the GW perturbations for a black-hole mimicker with a reflecting surface at $r_{s}=2M(1+\epsilon)$, with $R=1$ and $\epsilon=10^{-10}$. From the figure, it is clear that the graybody factor defined by Equation 38 approaches $1$ above $M\omega_{th}\gtrsim 0.39$, indicating transmission through the photon sphere. At low frequencies, we see sharp resonances associated with the real parts of the QNMs of the black hole mimicker, consistent with the plunge spectrum analysis in Sec. II.2.1 of the main text and the discussion in Appendix C.
where $dE_{22}/d\omega$ is our numerically computed energy flux from the main text, where the contribution to the source post-$t_{s}$ was ignored. From the numerical result presented in Sec. III, it can be seen that $(1/\mu^{2})\,dE_{22}/d\omega\sim 10^{-1}$ when $M\omega\simeq 0.7$ for all presented values of $\mathcal{L}$. So we find
$\frac{\mathcal{C}_{22\omega}}{\mu}\simeq 2.32\,M.$ (34)
We wish to now use Equation 31 to estimate the energy flux if we were to include the post-$t_{s}$ source term at $M\omega\simeq 0.7$. For this, we will first plug in Equations 28, 29 and 34 into the definition of $\alpha$ from Equation 32. So, for $(\ell,m)=(2,2)$, we have
$\alpha\sim\frac{|\delta\mathcal{C}_{22\omega}|}{\mathcal{C}_{22\omega}}\sim\frac{\tilde{\eta}\,T_{\rm post}}{2.32\,M},\qquad\tilde{\eta}\equiv\eta\kappa=\mathcal{O}(1),$ (35)
In particular, taking $\tilde{\eta}\sim 1$ and $T_{\rm post}\sim M$ yields $\alpha\simeq 0.43$.
Thus, we can get an estimate of the energy flux at $M\omega\simeq 0.7$ after accounting for the truncation effect using Equation 31, giving
[…]
Thus, under the minimal assumption that the post-$t_{s}$ source persists for a duration $T_{\rm post}=\mathcal{O}(M)$, the truncation-induced uncertainty in the spectrum at $M\omega\simeq 0.7$ (cf. Sec. III) is bounded to be at most a factor of a few; see Equation 36. Consequently, by referring to Sec. III, even after accounting for these truncation effects, the energy spectrum of the black-hole mimicker at $M\omega\simeq 0.7$ remains orders of magnitude larger than the corresponding Schwarzschild black-hole spectrum at the same $M\omega$.
[…]
Appendix C Graybody factors of a black hole mimicker and the energy spectrum.
In black hole perturbation theory, the greybody factor provides a quantitative measure of how efficiently perturbations transmit through the photon sphere, i.e. the peak of the effective Regge-Wheeler/Zerilli potential.
Graybody factors are defined from the homogeneous perturbations on the background spacetime. For each mode $(\ell,m)$ and frequency $\omega$, the radial master equation is a one-dimensional Schrödinger-like equation (cf. Equation 1), and homogeneous solutions at asymptotic infinity admit a standard scattering decomposition. For instance let us concentrate on the homogenous solution $X_{\ell m\omega}^{(\pm),\mathrm{in}}$ defined by Equation 8, with
$X_{\ell m\omega}^{(\pm),\mathrm{in}}\xrightarrow[r_{*}\to+\infty]{}\;A^{(\pm){\rm,in}}_{\ell m\omega}e^{-i\omega r_{*}}+A^{(\pm){\rm,out}}_{\ell m\omega}e^{+i\omega r_{*}},$ (37)
where $A^{\rm in/out}_{\ell m\omega}$ are the ingoing/outgoing amplitudes at asymptotic infinity. The potential barrier induces partial reflection of such perturbations, characterised by a notion of transmittivity also referred to as the greybody factor,
$\Gamma_{\ell m}(\omega)=\frac{1}{|A^{(\pm){\rm in}}_{\ell m\omega}|^{2}}.$ (38)
For a black hole mimicker, the above-defined quantity should be interpreted as a photon-sphere transmissivity, not a true absorption probability (more details regarding the subtleties in the notion of graybody factors for black hole mimickers can be found in Rosato:2025byu ).
Now, if for $\omega>\omega_{th}$ the associated perturbations are able to probe the near-horizon region more effectively, we expect the graybody factor defined through Equation 38 to approach $1$ when $\omega\gtrsim\omega_{th}$, in fact, this is precisely what we observe in Order of magnitude estimate of truncation-induced corrections at a representative frequency
We further note that this notion of graybody factor is not merely a convenient scattering diagnostic; it also enters directly as an overall prefactor in the radiated energy flux measured at asymptotic infinity and thus will carry the same low-frequency resonance comb in $dE/d\omega$ as discussed in the main body. To see this one can use Equations 11 and 12 along with the expression for the Wronskian between $X_{\ell m\omega}^{(\pm),\mathrm{in}}$ and $X_{\ell m\omega}^{(\pm),\mathrm{up}}$ from Equation 9 to obtain $dE/d\omega\sim\Gamma_{\ell m}(\omega)\times{\rm(inegral\,over\,source)}$. Thus, we can analytically expect the low-frequency resonance comb associated with the mimicker QNM spectrum demonstrated in Sec. II.2.1 to appear in the graybody factor plot as well, consistent with Order of magnitude estimate of truncation-induced corrections at a representative frequency.